# QNM families: classification and competition

### v1

@@ IV Complex amplitude and energy fraction
Thus, it is significant to develop the windowed analysis of the ringdown waveform. To start with, let us briefly review the standard procedure of fitting ringdown waveform in a data-driven framework. We shall firstly solve the master equation (3) using the finite difference method (see Appendix E.1 for details) to obtain the time-domain profile of the master scalar $\Psi(t)$. Second, we model the resulting time-series data based on the linear superposition of damped exponents ansatz, $\Psi(t)\simeq\sum_{k=1}^{p}h_{k}\,e^{-\mathrm{i}\omega_{k}t}$, where the sum over $p$ includes all relevant modes $\{\omega_{n}\}$ and their mirror modes $\{-\overline{\omega_{n}}\}$. Assuming that the modeling ringing phase starts at the peak of the signal $t=t_{\text{peak}}=0$ and ends at $t=N\,\Delta t_{\text{eff}}$, where $\Delta t_{\text{eff}}$ is the sampling time step and $N\geq 2p-1$ is an integer, we can write the discrete time-series signal $\Psi(m\,\Delta t_{\text{eff}})\equiv x[m]$ as
$x[m]=\sum_{k=1}^{p}h_{k}(z_{k})^{m}\,,\quad\text{with}\quad z_{k}=e^{-\mathrm{i}\omega_{k}\Delta t_{\text{eff}}}\,$, (12)
where $h_{k}$ is the complex amplitude of the $k$-th mode for the master scalar $\Psi(t)$. To connect this amplitude to a physical observable, we note that the outgoing EM radiation at infinity is described by the Maxwell-Newman-Penrose scalar $\mathbf{\phi}_{2}$ Teukolsky:1973ha , and as demonstrated by Chandrasekhar Chandrasekhar:1985kt , the RWZ master scalar $\Psi$ and the $\mathbf{\phi}_{2}$ are linked by a well-established transformation involving a first-order differential operator. For any given mode $k$ oscillating as $e^{-\mathrm{i}\omega_{k}t}$, this relationship manifests in the frequency domain as a direct proportionality to the mode’s frequency. Therefore, the complex amplitude $h_{k}$ of the master scalar can be related to the QNM excitation coefficient $C_{k}$ Berti:2025hly . As addressed in Andersson:1996cm ; Leaver:1986gd ; Berti:2006wq ; Oshita:2021iyn , the QNM excitation coefficient $C_{k}=E_{k}T_{k}$ can be further factorized into a source-independent excitation factor $E_{k}$ and a source-dependent term $T_{k}$. The factor $E_{k}$ is determined by the residue of the Green’s function at the corresponding pole, while $T_{k}$ depends on the initial data and on the specific source exciting the perturbations Andersson:1996cm . More studies on the QNM excitation coefficient can also be seen in Oshita:2025ibu ; Oshita:2024wgt ; Lo:2025njp and references therein. Here we focus on the complex amplitude $h_{k}$ for our following purpose. As suggested in Berti:2007dg , one can robustly extract $\{\omega_{k}\}$ via applying the MP method and subsequently determine $\{h_{k}\}$ using a linear least-squares fit, the main steps of which are present in Appendix E.2 and Appendix E.3, respectively.
@@ IV.1 Defining energy fraction in a sliding observation window
$\tilde{h}_{k}(t_{0})\equiv h_{k}(z_{k})^{d}=h_{k}e^{-\mathrm{i}\omega_{k}t_{0}}\,$, (16)
and its magnitude satisfies $|\tilde{h}_{k}|=|h_{k}|e^{-|\text{Im}(\omega_{k})|t_{0}}$, which encodes the decay of the mode up to the start of our observation. Therefore, the energy radiated by a single QNM in an observation window which is proportional to the time-integral of the square of the field’s time derivative Berti:2006hb , can be reduced as
$E_{k}(t_{0},T)$ $\propto$ $\int_{t_{0}}^{T+t_{0}}\left|\frac{\mathrm{d}\Psi_{k}(t)}{\mathrm{d}t}\right|^{2}\mathrm{d}t=\int_{t_{0}}^{T+t_{0}}|h_{k}(-\mathrm{i}\omega_{k})e^{-\mathrm{i}\omega_{k}t}|^{2}\mathrm{d}t$ (17)
$=$ $\int_{0}^{T}|\tilde{h}_{k}(t_{0})(-\mathrm{i}\omega_{k})e^{-\mathrm{i}\omega_{k}t^{\prime}}|^{2}\mathrm{d}t^{\prime}=|\tilde{h}_{k}(t_{0})|^{2}|\omega_{k}|^{2}\int_{0}^{T}e^{-2|\text{Im}(\omega_{k})|t^{\prime}}\mathrm{d}t^{\prime}$
$=$ $|\tilde{h}_{k}(t_{0})|^{2}|\text{Re}(\omega_{k})|^{2}\left(1+\frac{1}{4Q_{k}^{2}}\right)\int_{0}^{T}e^{-2|\text{Im}(\omega_{k})|t^{\prime}}\mathrm{d}t^{\prime}\approx|\tilde{h}_{k}(t_{0})|^{2}|\text{Re}(\omega_{k})|^{2}\int_{0}^{T}e^{-2|\text{Im}(\omega_{k})|t^{\prime}}\mathrm{d}t^{\prime},$
@@
$E_{k}(t_{0},T)\propto\frac{|\tilde{h}_{k}(t_{0})|^{2}|\text{Re}(\omega_{k})|^{2}}{2|\text{Im}(\omega_{k})|}\Big(1-e^{-2|\text{Im}(\omega_{k})|T}\Big)\,.$ (18)
@@
$\mathcal{E}_{k}(t_{0},T)\equiv\frac{E_{k}(t_{0},T)}{\sum_{j}E_{j}(t_{0},T)}=\frac{|\tilde{h}_{k}(t_{0})|^{2}|\text{Re}(\omega_{k})|^{2}|\text{Im}(\omega_{k})|^{-1}(1-e^{-2|\text{Im}(\omega_{k})|T})}{\sum_{j}|\tilde{h}_{j}(t_{0})|^{2}|\text{Re}(\omega_{j})|^{2}|\text{Im}(\omega_{j})|^{-1}(1-e^{-2|\text{Im}(\omega_{j})|T})}\,$, (19)
@@
To close this subsection, we summarize our windowed analysis workflow of the ringdown waveform as
@@ IV.2 Competition for dominant modes
    Inner-Peak Dominant Case: The profile of effective potential in this scenario corresponds to the last second panel of Fig. 1 and the inset of 5 in tortoise coordinates. The typical results in this case are shown in Fig. 5 and Tab. 3 for $\alpha=7.30$ and $l_{0}/M=0.940$.
@@
    Equal-Height Peaks Case: In this scenario, the two potential peaks are of comparable height (see the fourth panel of 1 and the inset of 6), which represents the most trapping mechanism and the perturbation spectrum is almost dominated by the echo family. This allows us to study the dynamic competition within the echo family itself. The results are shown in Fig. 6 and Tab. 4 with $\alpha=6.88$ and $l_{0}/M=0.940$. Comparing to the previous cases, the echo signal in Fig. 6 is explicit.
@@ Appendix A Hairy Schwarzschild black hole from gravitational decoupling
$\mathrm{d}s^{2}=-\Big(1-\frac{2\mathsf{M_{s}}}{r}\Big)e^{\alpha\,k(r)}\mathrm{d}t^{2}+\left(1-\frac{2\mathsf{M_{s}}}{r}\right)^{-1}e^{-\alpha\,k(r)}\mathrm{d}r^{2}+r^{2}\mathrm{d}\Omega^{2}\,$, (22)
@@
$\quad e^{\alpha\,k(r)}=1-\frac{\alpha\,c_{0}}{r}\left(1-\frac{2\mathsf{M_{s}}}{r}\right)^{-1}\penalty 10000\ \penalty 10000\ \text{with}\penalty 10000\ \penalty 10000\ c_{0}=\text{Const.}\,$, (23)
@@
where $h\equiv e^{\alpha\,k(r)}$ and the prime denotes the derivative with respect to $r$. One may consider the deformation should be trivial at $r=r_{h}$ and $r\gg\mathsf{M_{s}}$. Since the deformed black hole admits a proper horizon $r_{h}\sim 2\mathsf{M_{s}}$ and a asymptotically Schwarzschild geometric, thus a simple example emerges as
$\Theta^{\theta}{}_{\theta}=\frac{\alpha}{2\kappa\,\mathsf{M_{s}}^{2}r}e^{-r/\mathsf{M_{s}}}\left(r-2\mathsf{M_{s}}\right)\,.$ (30)
@@
$h(r)=c_{1}-\alpha\frac{c_{2}-r\,e^{-r/\mathsf{M_{s}}}}{r-2\mathsf{M_{s}}}=1-\left(\frac{l_{0}}{r}-\alpha\,e^{-\frac{r}{\mathsf{M_{s}}}}\right)\left(1-\frac{2\mathsf{M_{s}}}{r}\right)^{-1},$ (31)
@@
$\text{with}\quad A(r)=B(r)=1-\frac{2M}{r}+\alpha\,e^{-\frac{r}{M-l_{0}/2}}\equiv f(r)\quad\text{and}\quad C(r)=r^{2}\,$, (32)
@@ Appendix C Hyperboloidal framework of hairy Schwarzschild black hole
Having the above first-order-in-time PDE system (49), and performing the Fourier transformation in $\tau$ with $u(\tau,\sigma)\sim u(\sigma)e^{\mathrm{i}\omega\tau}$, we arrive at the eigenvalue problem $Lu=\mathrm{i}\omega u$. To numerically compute the QNM spectra, we discretize the differential operator $L$ using a Chebyshev-Lobatto collocation scheme (pseudospectral method), which yields its matrix representation $\mathbf{L}$. This discretization effectively reduces the original infinite-dimensional spectral problem to a finite-dimensional algebraic eigenvalue problem
@@ Appendix D Direct integration method
$\Psi_{\text{inf}}(r)=e^{\mathrm{i}\omega\,r_{\star}(r)}\sum^{\mathcal{I}}_{k=0}\mathfrak{p}_{k}\,r^{-k}.$ (54)

### v2

@@ IV Complex amplitude and energy fraction
Thus, it is significant to develop the windowed analysis of the ringdown waveform. To start with, let us briefly review the standard procedure of fitting ringdown waveform in a data-driven framework. We shall firstly solve the master equation (3) using the finite difference method (see Appendix E.1 for details) to obtain the time-domain profile of the master scalar $\Psi(t)$. Second, we model the resulting time-series data based on the linear superposition of damped exponents ansatz, $\Psi(t)\simeq\sum_{k=1}^{p}h_{k}\,\mathrm{e}^{-\mathrm{i}\omega_{k}t}$, where the sum over $p$ includes all relevant modes $\{\omega_{n}\}$ and their mirror modes $\{-\overline{\omega_{n}}\}$. Assuming that the modeling ringing phase starts at the peak of the signal $t=t_{\text{peak}}=0$ and ends at $t=N\,\Delta t_{\text{eff}}$, where $\Delta t_{\text{eff}}$ is the sampling time step and $N\geq 2p-1$ is an integer, we can write the discrete time-series signal $\Psi(m\,\Delta t_{\text{eff}})\equiv x[m]$ as
$x[m]=\sum_{k=1}^{p}h_{k}(z_{k})^{m}\,,\quad\text{with}\quad z_{k}=\mathrm{e}^{-\mathrm{i}\omega_{k}\Delta t_{\text{eff}}}\,$, (12)
where $h_{k}$ is the complex amplitude of the $k$-th mode for the master scalar $\Psi(t)$. To connect this amplitude to a physical observable, we note that the outgoing EM radiation at infinity is described by the Maxwell-Newman-Penrose scalar $\mathbf{\phi}_{2}$ Teukolsky:1973ha , and as demonstrated by Chandrasekhar Chandrasekhar:1985kt , the RWZ master scalar $\Psi$ and the $\mathbf{\phi}_{2}$ are linked by a well-established transformation involving a first-order differential operator. For any given mode $k$ oscillating as $\mathrm{e}^{-\mathrm{i}\omega_{k}t}$, this relationship manifests in the frequency domain as a direct proportionality to the mode’s frequency. Therefore, the complex amplitude $h_{k}$ of the master scalar can be related to the QNM excitation coefficient $C_{k}$ Berti:2025hly . As addressed in early foundational works Andersson:1996cm ; Leaver:1986gd ; Berti:2006wq ; Oshita:2021iyn , the QNM excitation coefficient $C_{k}=E_{k}T_{k}$ can be further factorized into a source-independent excitation factor $E_{k}$ and a source-dependent term $T_{k}$. The factor $E_{k}$ is determined by the residue of the Green’s function at the corresponding pole, while $T_{k}$ depends on the initial data and on the specific source exciting the perturbations Andersson:1996cm . This inspired more studies on the QNM excitation coefficient, see for examples Oshita:2025ibu ; Oshita:2024wgt ; Lo:2025njp and references therein. More recently, a definitive, cross-validated dataset for the Kerr excitation factors has been established Motohashi:2024fwt ; Kubota:2025hjk , in which the authors claimed to resolve ambiguities present in the earlier literature. Here we focus on the complex amplitude $h_{k}$ for our following purpose. As suggested in Berti:2007dg , one can robustly extract $\{\omega_{k}\}$ via applying the MP method and subsequently determine $\{h_{k}\}$ using a linear least-squares fit, the main steps of which are present in Appendix E.2 and Appendix E.3, respectively.
@@ IV.1 Defining energy fraction in a sliding observation window
$\tilde{h}_{k}(t_{0})\equiv h_{k}(z_{k})^{d}=h_{k}\mathrm{e}^{-\mathrm{i}\omega_{k}t_{0}}\,$, (16)
and its magnitude satisfies $|\tilde{h}_{k}|=|h_{k}|\mathrm{e}^{-|\text{Im}(\omega_{k})|t_{0}}$, which encodes the decay of the mode up to the start of our observation. Therefore, the energy radiated by a single QNM in an observation window which is proportional to the time-integral of the square of the field’s time derivative Berti:2006hb , can be reduced as
$E_{k}(t_{0},T)$ $\propto$ $\int_{t_{0}}^{T+t_{0}}\left|\frac{\mathrm{d}\Psi_{k}(t)}{\mathrm{d}t}\right|^{2}\mathrm{d}t=\int_{t_{0}}^{T+t_{0}}|h_{k}(-\mathrm{i}\omega_{k})\mathrm{e}^{-\mathrm{i}\omega_{k}t}|^{2}\mathrm{d}t$ (17)
$=$ $\int_{0}^{T}|\tilde{h}_{k}(t_{0})(-\mathrm{i}\omega_{k})\mathrm{e}^{-\mathrm{i}\omega_{k}t^{\prime}}|^{2}\mathrm{d}t^{\prime}=|\tilde{h}_{k}(t_{0})|^{2}|\omega_{k}|^{2}\int_{0}^{T}\mathrm{e}^{-2|\text{Im}(\omega_{k})|t^{\prime}}\mathrm{d}t^{\prime}$
$=$ $|\tilde{h}_{k}(t_{0})|^{2}|\text{Re}(\omega_{k})|^{2}\left(1+\frac{1}{4Q_{k}^{2}}\right)\int_{0}^{T}\mathrm{e}^{-2|\text{Im}(\omega_{k})|t^{\prime}}\mathrm{d}t^{\prime}\approx|\tilde{h}_{k}(t_{0})|^{2}|\text{Re}(\omega_{k})|^{2}\int_{0}^{T}\mathrm{e}^{-2|\text{Im}(\omega_{k})|t^{\prime}}\mathrm{d}t^{\prime},$
@@
$E_{k}(t_{0},T)\propto\frac{|\tilde{h}_{k}(t_{0})|^{2}|\text{Re}(\omega_{k})|^{2}}{2|\text{Im}(\omega_{k})|}\Big(1-\mathrm{e}^{-2|\text{Im}(\omega_{k})|T}\Big)\,.$ (18)
@@
$\mathcal{E}_{k}(t_{0},T)\equiv\frac{E_{k}(t_{0},T)}{\sum_{j}E_{j}(t_{0},T)}=\frac{|\tilde{h}_{k}(t_{0})|^{2}|\text{Re}(\omega_{k})|^{2}|\text{Im}(\omega_{k})|^{-1}(1-\mathrm{e}^{-2|\text{Im}(\omega_{k})|T})}{\sum_{j}|\tilde{h}_{j}(t_{0})|^{2}|\text{Re}(\omega_{j})|^{2}|\text{Im}(\omega_{j})|^{-1}(1-\mathrm{e}^{-2|\text{Im}(\omega_{j})|T})}\,$, (19)
@@
To close this subsection, we summarize our windowed analysis workflow of the time-domain waveform as
@@ IV.2 Competition for dominant modes
    Inner-Peak Dominant Case: The profile of effective potential in this scenario corresponds to the last second panel of Fig. 1 and the inset of Fig. 5 in tortoise coordinates. The typical results in this case are shown in Fig. 5 and Tab. 3 for $\alpha=7.30$ and $l_{0}/M=0.940$.
@@
    Equal-Height Peaks Case: In this scenario, the two potential peaks are of comparable height (see the fourth panel of Fig. 1 and the inset of Fig. 6), which represents the most trapping mechanism and the perturbation spectrum is almost dominated by the echo family. This allows us to study the dynamic competition within the echo family itself. The results are shown in Fig. 6 and Tab. 4 with $\alpha=6.88$ and $l_{0}/M=0.940$. Comparing to the previous cases, the echo signal in Fig. 6 is explicit.
@@ Acknowledgement
Data & code Availability: Data and code to obtain the data in this work are made publicly available on GitHub.
@@ Appendix A Hairy Schwarzschild black hole from gravitational decoupling
$\mathrm{d}s^{2}=-\Big(1-\frac{2\mathsf{M_{s}}}{r}\Big)\mathrm{e}^{\alpha\,k(r)}\mathrm{d}t^{2}+\left(1-\frac{2\mathsf{M_{s}}}{r}\right)^{-1}\mathrm{e}^{-\alpha\,k(r)}\mathrm{d}r^{2}+r^{2}\mathrm{d}\Omega^{2}\,$, (22)
@@
$\quad\mathrm{e}^{\alpha\,k(r)}=1-\frac{\alpha\,c_{0}}{r}\left(1-\frac{2\mathsf{M_{s}}}{r}\right)^{-1}\penalty 10000\ \penalty 10000\ \text{with}\penalty 10000\ \penalty 10000\ c_{0}=\text{Const.}\,$, (23)
@@
where $h\equiv\mathrm{e}^{\alpha\,k(r)}$ and the prime denotes the derivative with respect to $r$. One may consider the deformation should be trivial at $r=r_{h}$ and $r\gg\mathsf{M_{s}}$. Since the deformed black hole admits a proper horizon $r_{h}\sim 2\mathsf{M_{s}}$ and a asymptotically Schwarzschild geometric, thus a simple example emerges as
$\Theta^{\theta}{}_{\theta}=\frac{\alpha}{2\kappa\,\mathsf{M_{s}}^{2}r}\mathrm{e}^{-r/\mathsf{M_{s}}}\left(r-2\mathsf{M_{s}}\right)\,.$ (30)
@@
$h(r)=c_{1}-\alpha\frac{c_{2}-r\,\mathrm{e}^{-r/\mathsf{M_{s}}}}{r-2\mathsf{M_{s}}}=1-\left(\frac{l_{0}}{r}-\alpha\,\mathrm{e}^{-\frac{r}{\mathsf{M_{s}}}}\right)\left(1-\frac{2\mathsf{M_{s}}}{r}\right)^{-1},$ (31)
@@
$\text{with}\quad A(r)=B(r)=1-\frac{2M}{r}+\alpha\,\mathrm{e}^{-\frac{r}{M-l_{0}/2}}\equiv f(r)\quad\text{and}\quad C(r)=r^{2}\,$, (32)
@@ Appendix C Hyperboloidal framework of hairy Schwarzschild black hole
Having the above first-order-in-time PDE system (49), and performing the Fourier transformation in $\tau$ with $u(\tau,\sigma)\sim u(\sigma)\mathrm{e}^{\mathrm{i}\omega\tau}$, we arrive at the eigenvalue problem $Lu=\mathrm{i}\omega u$. To numerically compute the QNM spectra, we discretize the differential operator $L$ using a Chebyshev-Lobatto collocation scheme (pseudospectral method), which yields its matrix representation $\mathbf{L}$. This discretization effectively reduces the original infinite-dimensional spectral problem to a finite-dimensional algebraic eigenvalue problem
@@ Appendix D Direct integration method
$\Psi_{\text{inf}}(r)=\mathrm{e}^{\mathrm{i}\omega\,r_{\star}(r)}\sum^{\mathcal{I}}_{k=0}\mathfrak{p}_{k}\,r^{-k}.$ (54)